Required information [The following information applies to the questions displayed below.] Donnie Hilfiger has two classes ofRequired: 1. Record each of these transactions. (If no entry is required for a particular transaction/event, select No JournRequired: 1. Record each of these transactions. (If no entry is required for a particular transaction/event, select No JournRequired: 1. Record each of these transactions. (If no entry is required for a particular transaction/event, select No JournRequired: 1. Record each of these transactions. (If no entry is required for a particular transaction/event, select No JournView transaction list Journal entry worksheet Record the declaration of a cash dividend on both common and preferred stock ofRequired: 1. Record each of these transactions. (If no entry is required for a particular transaction/event, select No JournRequired: 1. Record each of these transactions. (If no entry is required for a particular transaction/event, select No Journ
Answer #1
Debit Credit
March 01 Cash 39600 =1100*36
    Common Stock 11 =1100*0.01
    Additional Paid in Capital 39589
May 15 Treasury Stock 11600 =400*29
     Cash 11600
July 10 Cash 6800 =200*34
      Treasury Stock 5800 =200*29
    Additional Paid in Capital 1000
October 15 Cash 7800 =200*39
    Preferred Stock 200 =200*1
    Additional Paid in Capital 7600
December 01 Dividends 4560 =(300+3400+1100-400+200+200)*0.95
    Dividends payable 4560
December 15 NO Journal entry required
December 31 Dividends payable 4560
      Cash 4560